HOB+0055 ; disk_pos 6408.5 GB
This is a 8 TB module and will need a new module in place. Contacted Imogen who confirmed that she will arrange for a fresh module in place. (vk)
01:37 UT - Restart observations. Missed scans: 073-0133 to 073-0139b. Recording was not stopped though. (vk)
HOB+0094 ; disk_pos 81.0 GB
HOB+0098 ; disk_pos 0.0 GB
Prior warnings that cyclone is due at Yg at 00:00 UT. However, camera and radar indicate nothing more severe than overcast skies at experiment start time.